OpenStudy (anonymous):

how do I solve the equation e^ { 2 x } - 7 e^ { x } +10 = 0?

OpenStudy (zzr0ck3r):

let a = e^x then you have a^2-7a+10 = 0 (a-2)(a-5)=0 so e^x-2 = 0 and e^x-5=0 e^x=2 and e^x = 5 can you do it now?

OpenStudy (anonymous):

yes. thanks!

OpenStudy (zzr0ck3r):

np